2 The frst tive vibrational energy levels of HCl n wavenumbers are 1481 86 спı | 4367 50 cin_1 | 7149 04 cm 9826 48 m 12399 8 cm CImm Calculate the dissociation cnergies Do and De 1141
Answer #1

The vibrational energy levels given are,

/) = 148 186cm1-1. v1 = 436750cm-l. v-= 719404cm-1,V3 = 982648cm-i,

The required dissociation energies are given as,

D_0 = \nu_n - \nu_1 = \nu _3 - \nu _1 = 545898cm^-1

D_e = \nu_n - \nu_0 = \nu _3 - \nu _0 = 832462 cm^-1

Converting these vibrational energy levels in cm-1 to energy levels in joules,

D_0 = \frac{hc(545898)}{N_A} J/mol= 1801.6537 J/mol = 1.8 kJ/mol

D_e = \frac{hc(832462)}{N_A} J/mol= 2747.4148 J/mol = 2.75 kJ/mol

P.S. The fifth vibrational energy should be higher than the fourth, the value given was smaller and not considered. The calculation will follow the same pattern, i.e. D_0 = \nu_n - \nu_1 , D_e = \nu_n - \nu_0